OpenStudy (anonymous):

joan can mow a 10-acre field in 5 hours. How long would it take her to moa a 4-acre field?

OpenStudy (anonymous):

5/10 = x/4 20 = 10x x = 2

OpenStudy (anonymous):

10/5 = 4/x

OpenStudy (anonymous):

20 = 10x

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

or in your head. ten for 5 hours, 2 an hour. so 4 takes 2 hours